OpenStudy (bluemoon1):

Write the slope intercept form of the equation of the line through the given point with the given slope. Through (-3,3), slope = 1/3

OpenStudy (amorfide):

slope intercept form is y=mx+b m is the slope m=1/3 so write our the slope intercept form, by substituting m=1/3, then we will continue

OpenStudy (bluemoon1):

How do I substitute

OpenStudy (bluemoon1):

Y= (1/3)

OpenStudy (amorfide):

you replace m, with the value of m

OpenStudy (bluemoon1):

So is my answer above right?

OpenStudy (amorfide):